# Dark-to-black super accretion as a mechanism for early supermassive black hole growth

### v2

@@ Abstract
Introduction. Supermassive black holes (SMBHs), dark matter, and stars are the three fundamental components of galaxies. Their formation, growth, and evolution are highly coupled and require processes that in many cases last for hundreds to thousands of millions of years. However, recent observations reveal that massive galaxies hosting SMBHs were already in place at very high redshifts quasars at $z\sim 7$ for example are known to contain black holes of mass $\sim 10^{9}\,M_{\odot}$ Maiolino et al. (2024); Bañados et al. (2018), and very recently galaxy candidates have been identified out to $z\sim 16$ or more by the deep James Webb Space Telescope (JWST) surveys Atek et al. (2023); Robertson et al. (2024), suggesting significant structure formation when the Universe was less than a few hundred million years old. These findings strain the standard pathways by which black hole seeds grow via baryonic accretion and mergers within hierarchical structure formation Valiante et al. (2016).
@@
In parallel, bosonic fields (scalar, vector or axion-like) can resonate with black holes at different mass scales Arvanitaki and Dubovsky (2011); Herdeiro (2023). One of the best-known resonant phenomena is superradiance  Brito et al. (2015). Superradiance occurs for bosonic fields around rotating black holes, with modes with positive azimuthal number $m>0$ whose frequency lies below the critical threshold $\omega<m\Omega_{H}$, where $\Omega_{H}$ is the angular velocity of the horizon. Even then, the corresponding growth timescales are generally extremely long, rendering the effect negligible in most regimes. A notable exception arises when the dimensionless parameter $\mu M_{\rm BH}\sim\mathcal{O}(1)$, where $M_{\rm BH}$ is the black hole mass and $\mu$ is the scalar particle mass (in natural units, $c=G=\hbar=1$), for which the instability is significantly enhanced (see Dolan (2013); Herdeiro and Radu (2014); East and Pretorius (2017); Conlon and Herdeiro (2018); Degollado et al. (2018)). In this case, there exists a specific black hole mass range where superradiance becomes the dominant process shaping the dynamics of the scalar cloud.
In this paper we propose a related resonant mechanism, but instead of extracting energy via superradiance, the scalar field forms quasi-bound states around the black hole which decay exponentially, i.e. an effective superaccretion of the bosonic field. Massive bosonic dark matter is known to support quasi-bound states in the vicinity of black holes Detweiler (1980a); Barranco et al. (2011, 2012); Sanchis-Gual et al. (2015a, b, 2016), whose decay timescales depend sensitively on both the black hole mass and the field mass. The accretion rate depends on whether the Compton wavelength of the boson $\lambda_{c}\sim\frac{1}{\mu}$ is greater than the black hole horizon scale $r_{\rm H}$, if $\lambda_{c}>>r_{\rm H}$, the boson’s wave function spreads out and the black hole only absorbs a very small fraction. In this case the cross-section is much smaller than the geometric area of the horizon and the accretion rate is suppressed. For very small $\mu$ (e.g. $\mu\sim 10^{-22}$$10^{-20}$ eV) and large SMBHs, the accretion (or absorption) rate is negligible over cosmological timescales Urena-Lopez and Liddle (2002); Annulli et al. (2020); however, depending on the particle mass and the black hole growth history, there can be regimes where the timescale becomes comparable to or shorter than cosmic ages.
@@
In what follows, using approximations adapted from the quasi-bound state calculations Barranco et al. (2011, 2012, 2014), we integrate the accretion rate of scalar field dark matter over cosmological time as a function of $\mu$ and black hole mass. We demonstrate that this dark-to-black conversion leads to exponential black hole growth. Modest seeds in the range $10^{2}$$10^{5}M_{\odot}$ embedded in bosonic clouds of $\sim 10^{6}-10^{8}M_{\odot}$ can reach supermassive scales within $\sim 10^{8}$ yrs. This process operates independently of baryonic accretion, providing a natural and testable mechanism to explain the emergence of high-redshift SMBHs and linking early-universe black hole demographics to fundamental dark matter physics.
@@
Framework. We analyze the accretion of massive scalar fields into black holes within the framework of general relativity within the Einstein’s equations $R_{\alpha\beta}-\frac{1}{2}Rg_{\alpha\beta}=8\pi T_{\alpha\beta}$, where $R_{\alpha\beta}$ and $R$ are the Ricci tensor and Ricci scalar associated to the metric $g_{\alpha\beta}$. To isolate the essential features of the accretion process, we work in the test-field limit, neglecting the backreaction of the scalar field on the spacetime geometry. The background is assumed to be spherically symmetric, with the line element taken as $ds^{2}=-f(r)dt^{2}+\frac{1}{f(r)}dr^{2}+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\varphi^{2}\right),$ where $f(r)=1-\frac{r_{H}}{r}$ with $r_{H}=2M_{\rm BH}$. Conservation of the stress energy tensor implies that the scalar field obeys the Klein Gordon equation $\nabla_{\alpha}\nabla^{\alpha}\Psi=\mu^{2}\Psi$. We focus on quasi-stationary states, for which the scalar field has the form $\Psi=e^{-i\sigma t}\psi(r)$. Furthermore, we are interested in solutions of the Klein–Gordon equation, subject to ingoing boundary conditions at the horizon and outgoing behavior at spatial infinity. The spectral properties of this system have been extensively studied Damour et al. (1976); Zouros and Eardley (1979); Detweiler (1980b); Furuhashi and Nambu (2004). In the regime $\mu M_{BH}<<1$, the quasi–bound state spectrum of the massive scalar field closely resembles that of the hydrogen atom with a complex frequency $\sigma$ whose real ($\omega$) and imaginary ($\Gamma$) parts are given, for a spherically symmetric distribution, as:
$M_{\rm BH}\omega$ $=$ $\mu M_{\rm BH}\left[1-(\mu M_{\rm BH})^{2}\right]^{1/2}\;$
$M_{\rm BH}\Gamma$ $=$ $-16(\mu M_{\rm BH})^{6}\;.$ (1)
@@
Accretion of matter and black hole growth. To model the mass growth of the black hole, we adopt an adiabatic approximation in which the scalar field is gradually absorbed by the horizon, leading to a corresponding increase in the black hole mass. The black hole evolves smoothly, while at each instant the scalar field can be regarded as approximately stationary with respect to the background geometry. The black hole mass $M_{\rm BH}(t)$ grows at the expense of the cloud mass $M_{c}(t)$ according to
@@
where the accretion rate is given by Eq. (1). The equations conserve the total mass $M_{\rm BH}+M_{c}$ and yield in the late stages an exponential depletion of the cloud. We solve the system using a high accuracy explicit Runge–Kutta method (5th order with embedded 4th order) with adaptive step size, terminating the integration when the cloud mass falls below $10^{-15}M_{\odot}$. We extract the instantaneous accretion rate $\dot{M}_{\rm BH}$ in $M_{\odot}\,{\rm yr}^{-1}$.
@@
In Fig. 3 we present the evolution of the accretion rate and of the dimensionless parameter $\mu M_{\rm BH}$, offering a different perspective on the same dynamics. As the system evolves, $\mu M_{\rm BH}$ increases steadily and saturates at a final value of $\sim 0.13$. This illustrates how the black hole mass growth naturally drives the system into the resonant regime where accretion becomes most efficient in the expected timescale. Taken together with the runaway behavior discussed above, these results demonstrate a possible pathway for black holes to experience accelerated growth phases, providing a natural mechanism to account for the emergence of supermassive black holes already observed at high redshift.
@@
Discussion and Outlook. The process described here can be thought of as a dark-to-black conversion: dark matter becomes part of the black holes, losing casual contact with the rest of the Universe. This channel would have been particularly efficient in the early Universe, when proto-galaxies were strongly dark matter dominated and baryonic collapse was still inefficient. The selective nature of the resonance condition ($\mu M_{\rm BH}\sim 1$) implies that only a fraction of the scalar field component participates in the process, thus avoiding tensions with large-scale structure and Lyman-$\alpha$ forest constraints Iršič et al. (2017); Armengaud et al. (2017), while still producing a dynamically relevant effect, in good agreement with recent studies in the full non-linear regime Cardoso et al. (2022). This efficient transfer would simultaneously accelerate the growth of SMBHs and reduce the central dark matter density in their vicinity, deepening the local potential well and enhancing subsequent baryonic inflows. Its observational imprints would be indirect, manifesting deviations from standard halo profiles, lensing signatures, or SMBH–host scaling relations at high redshift. In this way, dark-to-black conversion could represent a key missing link between dark matter physics and the unexpectedly rapid emergence of massive black holes in the early Universe.
@@
Because the resonance is sensitive to $M_{\rm BH}$ and $\mu$ and the cloud supplies are finite, this channel is effectively a one-time conversion per halo, that is, once the local coherent reservoir is depleted by accretion, the same galaxy cannot generically repeat the event for smaller black holes formed later. Any remnant cloud would be small and the process leaves only gravitational imprints (modified central halo profiles, lensing anomalies, or altered SMBH–host scaling) or indirect probes via superradiance/spin limits in the appropriate mass windows Cardoso et al. (2018); Cunha et al. (2019). Although we have focused on minimally coupled scalar fields, analogous quasibound phenomena exist for other bosonic fields Rosa and Dolan (2012), massive Dirac fields Lasenby et al. (2005); Dolan and Dempsey (2015) and for mixed boson–fermion couplings. Recent work has revisited fermionic quasibound spectra around Schwarzschild/Kerr black holes and explored boson-driven fermion production, suggesting that the qualitative resonant absorption channel may extend beyond scalars to a broader class of dark-sector candidates.
Acknowledgements. NSG acknowledges support from the Spanish Ministry of Science and Innovation via the Ramón y Cajal programme (grant RYC2022-037424-I), funded by MCIN/AEI/10.13039/501100011033 and by “ESF Investing in your future”. This work is further supported by the Spanish Agencia Estatal de Investigación (Grant PID2021-125485NB-C21 and PID2024-159689NB-C21) funded by MCIN/AEI/10.13039/501100011033 and ERDF A way of making Europe, and by the European Horizon Europe staff exchange (SE) programme HORIZON-MSCA2021-SE-01 Grant No. NewFunFiCO-101086251. This work is also supported by CIDMA under the FCT Multi-Annual Financing Program for R&D Units (UID/4106/2025) through the Portuguese Foundation for Science and Technology (FCT), project 2024.05617.CERN (https://doi.org/10.54499/2024.05617.CERN) and by DGAPA-UNAM through grant IN110523.

### v3

@@ Abstract
I Introduction
Supermassive black holes (SMBHs), dark matter, and stars are the three fundamental components of galaxies. Their formation, growth, and evolution are highly coupled and require processes that in many cases last for hundreds to thousands of millions of years. However, recent observations reveal that massive galaxies hosting SMBHs were already in place at very high redshifts quasars at $z\sim 7$ for example are known to contain black holes of mass $\sim 10^{9}\,M_{\odot}$ Maiolino et al. (2024); Bañados et al. (2018), and very recently galaxy candidates have been identified out to $z\sim 16$ or more by the deep James Webb Space Telescope (JWST) surveys Atek et al. (2023); Robertson et al. (2024), suggesting significant structure formation when the Universe was less than a few hundred million years old. These findings strain the standard pathways by which black hole seeds grow via baryonic accretion and mergers within hierarchical structure formation Valiante et al. (2016).
@@ I Introduction
In parallel, bosonic fields (scalar, vector or axion-like) can resonate with black holes or even neutron stars at different mass scales Arvanitaki and Dubovsky (2011); Herdeiro (2023); Lazarte et al. (2025). One of the best-known resonant phenomena is superradiance  Brito et al. (2015). Superradiance occurs for bosonic fields around rotating black holes, for modes with positive azimuthal number $m>0$ whose frequency lies below the critical threshold $\omega<m\Omega_{H}$, where $\Omega_{H}$ is the angular velocity of the horizon. Even then, the corresponding growth timescales are generally extremely long, rendering the effect negligible in most regimes. A notable exception arises when the dimensionless parameter $\mu M_{\rm BH}\sim\mathcal{O}(1)$, where $M_{\rm BH}$ is the black hole mass and $\mu$ is the scalar particle mass (in natural units, $c=G=\hbar=1$), for which the instability is significantly enhanced (see Dolan (2013); Herdeiro and Radu (2014); East and Pretorius (2017); Conlon and Herdeiro (2018); Degollado et al. (2018)). In this case, there exists a specific black hole mass range where superradiance becomes the dominant process shaping the dynamics of the scalar cloud.
In this paper we propose a related resonant mechanism, but instead of extracting energy via superradiance, the scalar field forms quasi-bound states around the black hole which decay exponentially, i.e. an effective superaccretion of the bosonic field. Massive bosonic dark matter is known to support quasi-bound states in the vicinity of black holes Detweiler (1980a); Barranco et al. (2011, 2012); Sanchis-Gual et al. (2015a, b, 2016), whose decay timescales depend sensitively on both the black hole mass and the field mass. The accretion rate depends on whether the Compton wavelength of the boson $\lambda_{c}\sim\frac{1}{\mu}$ is greater than the black hole horizon scale $r_{\rm H}$. If $\lambda_{c}\gg r_{\rm H}$, the boson’s wave function spreads out and the black hole only absorbs a very small fraction. In this case the cross-section is much smaller than the geometric area of the horizon and the accretion rate is suppressed. For very small $\mu$ (e.g. $\mu\sim 10^{-22}$$10^{-20}$ eV) and large SMBHs, the accretion (or absorption) rate is negligible over cosmological timescales Urena-Lopez and Liddle (2002); Annulli et al. (2020); however, depending on the particle mass and the black hole growth history, there can be regimes where the timescale becomes comparable to or shorter than cosmic ages.
@@
In what follows, using approximations adapted from the quasi-bound state calculations Barranco et al. (2011, 2012, 2014a), we integrate the accretion rate of scalar field dark matter over cosmological time as a function of $\mu$ and black hole mass. We demonstrate that this dark-to-black conversion leads to exponential black hole growth. Modest seeds in the range $10^{2}$$10^{5}M_{\odot}$ embedded in bosonic clouds of $\sim 10^{6}-10^{8}M_{\odot}$ can reach supermassive scales within $\sim 10^{8}$ yrs. This process operates independently of baryonic accretion, providing a natural and testable mechanism to explain the emergence of high-redshift SMBHs and linking early-universe black hole demographics to fundamental dark matter physics.
II Setup
@@
We analyze the accretion of massive scalar fields into black holes within the framework of general relativity within Einstein’s equations $R_{\alpha\beta}-\frac{1}{2}Rg_{\alpha\beta}=8\pi T_{\alpha\beta}$, where $R_{\alpha\beta}$ and $R$ are the Ricci tensor and Ricci scalar associated to the metric $g_{\alpha\beta}$. To isolate the essential features of the accretion process, we work in the test-field limit, neglecting the backreaction of the scalar field on the spacetime geometry. The background is assumed to be spherically symmetric, with the line element taken as $ds^{2}=-f(r)dt^{2}+\frac{1}{f(r)}dr^{2}+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\varphi^{2}\right),$ where $f(r)=1-\frac{r_{H}}{r}$ with $r_{H}=2M_{\rm BH}$. Conservation of the stress energy tensor implies that the scalar field obeys the Klein Gordon equation $\nabla_{\alpha}\nabla^{\alpha}\Psi=\mu^{2}\Psi$. We focus on quasi-stationary states, for which the scalar field has the form $\Psi=e^{-i\sigma t}\psi(r)$. Furthermore, we are interested in solutions of the Klein–Gordon equation, subject to ingoing boundary conditions at the horizon and outgoing behavior at spatial infinity. The spectral properties of this system have been extensively studied Damour et al. (1976); Zouros and Eardley (1979); Detweiler (1980b); Furuhashi and Nambu (2004), and the behavior of scalar fields in the non-rotating case has been analyzed in detail Barranco et al. (2011, 2012, 2014a, 2014b). In the regime $\mu M_{\rm BH}\ll 1$, it has been shown that the dynamical accretion of generic scalar field configurations onto a Schwarzschild black hole is well described by the quasi–bound state spectrum of a massive scalar field Barranco et al. (2012); Baumann et al. (2019), concluding that such spectrum closely resembles that of the hydrogen atom, with a complex frequency $\sigma=\omega+i\gamma$, whose real ($\omega$) and imaginary ($\gamma$) parts are given, for a spherically symmetric distribution, by
$M_{\rm BH}\omega$ $=$ $\mu M_{\rm BH}\left[1-(\mu M_{\rm BH})^{2}\right]^{1/2},$
$M_{\rm BH}\gamma$ $=$ $-8(\mu M_{\rm BH})^{6},$ (1)
where the imaginary part $\gamma$ encodes the absorption of scalar field at the black hole horizon. Although these expressions are derived in the test-field approximation, we use them to model a slow, gradual mass transfer from the scalar cloud to the black hole, in which the accreted mass induces a small incremental change in $M_{\rm BH}$ and a corresponding depletion of the cloud, in a cyclic process. This approximation is supported by fully relativistic numerical simulations Cardoso et al. (2022). As shown in Refs. Detweiler (1980a); Baumann et al. (2019), the dominant contribution arises from the ground state ($n=0$ in our case), while higher-$n$ states are subdominant.
@@
The decay of the scalar field amplitude is characterized by $\gamma$, but we will instead consider the decay of the scalar field energy given by
$\Gamma=2\gamma=-16\mu^{6}M_{\rm BH}^{5}.$ (2)
@@
III Accretion of matter and black hole growth.
To model the mass growth of the black hole, we adopt an adiabatic approximation in which the scalar field is gradually absorbed by the horizon, leading to a corresponding increase in the black hole mass. The black hole evolves smoothly, while at each instant the scalar field can be regarded as approximately stationary with respect to the background geometry. The black hole mass $M_{\rm BH}(t)$ grows at the expense of the cloud mass $M_{c}(t)$ according to
@@
where the accretion rate is given by Eq. (1). The equations conserve the total mass $M_{\rm BH}+M_{c}$ and yield in the late stages an exponential depletion of the cloud. We assume that the cloud loses energy solely via absorption at the black hole horizon. In exact spherical symmetry there is no gravitational radiation, and scalar radiation is strongly suppressed during the rapid accretion phase, at least for complex scalar fields, as shown by fully relativistic simulations Cardoso et al. (2022). Relaxing spherical symmetry would generically allow additional loss channels and alter the effective accretion rate.
In our scenario, the scalar cloud represents a dense dark matter component surrounding the black hole. In fuzzy (ultralight scalar) dark matter cosmologies, solitonic cores naturally arise from gravitational collapse of small initial density fluctuations, forming on roughly the halo free-fall timescale, $t_{\rm cond}\sim 10^{8}$$10^{9}\,$yrs for particle masses $\mu\sim 10^{-22}$$10^{-20}\,$eV Schive et al. (2014); Hui et al. (2017); Davies and Mocz (2020); Hui (2021). Simulations show such cores exist at high redshift ($z\gtrsim 7$) with masses $M\gtrsim 10^{9}\,M_{\odot}$ Chiu et al. (2025), making a massive scalar cloud around an early black hole seed astrophysically plausible. For the scalar masses considered here ($\mu\sim 10^{-19}$$10^{-15}\,\mathrm{eV}$), the cloud should not be directly interpreted as a kiloparsec-scale halo or solitonic core. Instead, it can be modeled as a compact bosonic cloud overdensity formed through gravitational cooling Seidel and Suen (1994); Liebling and Palenzuela (2023); Di Giovanni et al. (2018), as commonly found in the nonlinear evolution of self-gravitating scalar fields. Such configurations have a characteristic size $R\sim\mu^{-1}$ and form on dynamical timescales $t\sim R\sim\mu^{-1}$, which can be potentially much shorter than the initial epoch considered in our analysis ($t_{i}\sim 10^{9}\,\mathrm{yrs}$). In this work, we therefore treat the scalar cloud as a pre-existing reservoir of dark matter surrounding the black hole and focus on its subsequent accretion.
We solve the system using a high accuracy explicit Runge–Kutta method (5th order with embedded 4th order) with adaptive step size, terminating the integration when the cloud mass falls below $10^{-15}M_{\odot}$. We extract the instantaneous accretion rate $\dot{M}_{\rm BH}$ in $M_{\odot}\,{\rm yr}^{-1}$.
@@
In Fig. 3 we present the evolution of the accretion rate $\dot{M}_{\rm BH}$ and of the dimensionless parameter $\mu M_{\rm BH}$, offering a different perspective on the same dynamics. As the system evolves, $\mu M_{\rm BH}$ increases steadily and saturates at a final value of $\sim 0.13$. In the top panel of Fig. 3, the black dashed horizontal line marks the Eddington accretion limit for the chosen black hole seed mass. Therefore, the accretion rate becomes super-Eddington for a short period of time. This illustrates how the black hole mass growth naturally drives the system into the resonant regime where accretion becomes most efficient in the expected timescale. Taken together with the runaway behavior discussed above, these results demonstrate a possible pathway for black holes to experience accelerated growth phases, providing a natural mechanism to account for the emergence of supermassive black holes already observed at high redshift.
@@
IV Discussion and Outlook
The process described here can be thought of as a dark-to-black conversion: dark matter becomes part of the black holes, losing casual contact with the rest of the Universe. This channel would have been particularly efficient in the early Universe, when proto-galaxies were strongly dark matter dominated and baryonic collapse was still inefficient. The selective nature of the resonance condition ($\mu M_{\rm BH}\sim 1$) implies that only a fraction of the scalar field component participates in the process, thus avoiding tensions with large-scale structure and Lyman-$\alpha$ forest constraints Iršič et al. (2017); Armengaud et al. (2017), while still producing a dynamically relevant effect, in good agreement with recent studies in the full non-linear regime Cardoso et al. (2022). This efficient transfer would simultaneously accelerate the growth of SMBHs and reduce the central dark matter density in their vicinity, deepening the local potential well and enhancing subsequent baryonic inflows. Its observational imprints would be indirect, manifesting deviations from standard halo profiles, lensing signatures, or SMBH–host scaling relations at high redshift. In this way, dark-to-black conversion could represent a key missing link between dark matter physics and the unexpectedly rapid emergence of massive black holes in the early Universe.
@@ IV Discussion and Outlook
Because the resonance is sensitive to $M_{\rm BH}$ and $\mu$ and the cloud supplies are finite, this channel is effectively a one-time conversion per halo, that is, once the local coherent reservoir is depleted by accretion, the same galaxy cannot generically repeat the event for smaller black holes formed later. Any remnant cloud would be small and the process leaves only gravitational imprints (modified central halo profiles, lensing anomalies, or altered SMBH–host scaling) or indirect probes via superradiance/spin limits in the appropriate mass windows Cardoso et al. (2018); Cunha et al. (2019). Although we have focused on minimally coupled scalar fields, analogous quasibound phenomena exist for other bosonic fields Rosa and Dolan (2012), massive Dirac fields Lasenby et al. (2005); Dolan and Dempsey (2015) and for mixed boson–fermion couplings. Recent works have revisited fermionic quasibound spectra around Schwarzschild/Kerr black holes and explored boson-driven fermion production Chen et al. (2025a, b), suggesting that the qualitative resonant absorption channel may extend beyond scalars to a broader class of dark-sector candidates.
Acknowledgements. NSG acknowledges support from the Spanish Ministry of Science, Innovation, and Universities via the Ramón y Cajal programme (grant RYC2022-037424-I), funded by MCIN/AEI/10.13039/501100011033 and by “ESF Investing in your future”. This work is further supported by the Spanish Agencia Estatal de Investigación (Grant PID2021-125485NB-C21 and PID2024-159689NB-C21) funded by MCIN/AEI/10.13039/501100011033 and ERDF A way of making Europe, and by the European Horizon Europe staff exchange (SE) programme HORIZON-MSCA2021-SE-01 Grant No. NewFunFiCO-101086251. This work is also supported by CIDMA under the FCT Multi-Annual Financing Program for R&D Units (UID/4106/2025) through the Portuguese Foundation for Science and Technology (FCT), project 2024.05617.CERN (https://doi.org/10.54499/2024.05617.CERN) and by DGAPA-UNAM through grant IN110523.